OpenStudy (anonymous):

Can someone help me solve, 4sin^2(theta)=3+cos^2(theta), for values of theta between 0 and 360 degrees. thank you :)

OpenStudy (anonymous):

sin^2 (t) +cos^2(t) = 1 (by definition) so, cos^2(t) = 1-sin^2(t) substitute: 4sin^2(t) =3+1-sin^2(t) add sin^2(t) to both sides: 5sin^2(t)=4 divide by 5 both sides: sin^2(t) = 4/5 sin(t)= 2/sq.rt(5)

OpenStudy (anonymous):

thanks :)